OpenStudy (anonymous):

What is the antiderivative for 1/x^2-2x+2 ?

OpenStudy (goformit100):

Just for it you have to that Integrate 1/x^2-2x+2.

OpenStudy (anonymous):

How do I do it?

OpenStudy (goformit100):

@abb0t Plz Help Us.

OpenStudy (abb0t):

To integrate: \(\int\limits a^ndx= \frac{ a^{n+1} }{ n+1 }+C\) where C is some arbitrary constant. Also, NOTE: \[\frac{ 1 }{ a^n } = 1a^{-n}\]

geerky42 (geerky42):

@Dharmaputra_993 (x-1)² = x²-2x+1

OpenStudy (anonymous):

1/(x^2-2x+2) = 1/((x - 1)^2 + 1), completing the square. Let x - 1 = tan(u), so that 1/((x - 1)^2 + 1) = 1/(tan^2(u) + 1) = 1/sec^2(u); and d(x - 1) = d(tan(u)) dx = sec^2(u)du; so the antiderivativ goes to integral(1/((x - 1)^2 + 1))dx = integral(sec^2(u)/sec^2(u))du = integral(du) = u + C and since x - 1 = tan(u), u = arctan(x - 1), so this is just arctan(x - 1) + C.

OpenStudy (anonymous):

Right, sorry.

OpenStudy (abb0t):

Oh, oops. Yeah, I think you want to use partial fractions to solve this.

geerky42 (geerky42):

Xxnina's approach is better.

OpenStudy (abb0t):

I misread the question, sorry about that.

OpenStudy (anonymous):

got it thank you (:

OpenStudy (anonymous):

x^2-2x+2= (x-1)^2+1 so antiderivate of 1/x^2-2x+2 = antiderivate of 1/((x-1)^2+1) = arctan (x-1) + C Good Luck! I hope you understand!

geerky42 (geerky42):

x² - 2x + 2 = (x-1)² + 1

OpenStudy (anonymous):

yeah i was stuck at how to approach it thank you so much. Nina did x^2-2x+2+1i believe.

OpenStudy (abb0t):

I think it's -arctan

OpenStudy (anonymous):

that would make it (1-x)

OpenStudy (anonymous):

yes! great work! :)

OpenStudy (abb0t):

(1-x) is the correct form of writing it usually. As you get higher in math, you'll notice that it's often writen as (1-x) rather than (-x+1) or (x-1). It has to do with series.